### v1

@@ A.1.1 Variational EM algorithm with MM updates
where $\bm{\alpha},\bm{\beta},\bm{\eta}$ and $\bm{\xi}^{(s)}$ are fixed. Maximizing the function $M$ guarantees that the lower bound does not decreases. Extending the approach of Vu et al. (2013) to directed networks, the following expression minorizes $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\xi})$
$M\left(\bm{\xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\xi}^{(s)}\right)$ ::=:= $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}(g_{ij},\bm{x}_{ij},\bm{z})$
@@
In Section A.1.2, we detail how sparse matrix multiplication can be exploited to speed up the updates of $\bm{\xi}^{(s+1)}$
We run this algorithm until the relative increase of the variational lower bound in below a certain threshold to obtain estimates for $\widehat{\bm{\xi}}$ and $\widehat{\bm{\eta}}$. We then assign each node to its modal estimated type, such that $\widehat{z}_{ik}=1$ if $\widehat{\bm{\xi}}_{ik}\geq\widehat{\bm{\xi}}_{i\ell}$ for all $\ell\neq k$ and for all $i$s.
A.1.2 Exploiting sparse matrix operations for updating variational parameters
Next, we detail how the updates of $\mathbf{\xi}$ in Step 1 can be carried out in a scalable manner based on sparse matrix operations. Equation (A.1.1) can be written as a quadratic programming problem in $\mathbf{\xi}$ with the side constraints that $\sum_{k=1}^{K}\xi_{i,k}=1$ for all $i=1,\ldots,N$ needs to hold. Therefore, we first rearrange the first part of (A.1.1):
$\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}(g_{ij},\bm{x}_{ij})$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\frac{\xi_{ik}^{2}}{2\xi_{ik}^{(s)}}\xi_{jl}^{(s)}\left\{\log\,\pi_{kl}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right\}$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\frac{\Omega_{ik}^{(s)}(\bm{g},\bm{x},\bm{\xi}^{(s)})}{2\xi_{ik}^{(s)}}\xi_{ik}^{2}$
@@ A.1.2 Exploiting sparse matrix operations for updating variational parameters
$\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)~{}\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left\{\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right\}.$ (16)
@@
$M\left(\bm{\xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\xi}^{(s)}\right)$ $=\sum_{i=1}^{N}\sum_{k=1}^{K}\left(\frac{\Omega_{ik}^{(s)}(\bm{g},\bm{x},\bm{\xi}^{(s)})}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}\right)\xi_{ik}^{2}+\left(\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1\right)\xi_{ik}$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}A_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)\,\xi_{ik}^{2}+B_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)\,\xi_{ik}$
@@
$A_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)\coloneqq\frac{\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}$
@@
$B_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)\coloneqq\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1$
@@
To update the estimate of $\bm{\xi}$, we need to evaluate $A_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ and $B_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$, which, when done naively, is of complexity $O(N^{2}\,K^{2})$. Computing $\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ for $A_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ is the leading term driving the algorithmic complexity. It is thus prohibitively difficult for a large number of population members and communities in the network. Note that in our application, this problem is exasperated as we have more than 35,000 nodes and our theoretical result on the adequacy of the estimation procedure assumes that $K$ grows as a function of $N$. To avoid this issue, we show how $\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ can be evaluated through a series of matrix multiplications. The underlying idea of our approach is that $\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ has an easy form if $\mathbf{g}$ is an empty network with all pairwise covariates set to zero, i.e., $g_{ij}=0$ and $x_{q}=0\,\forall\,i,j=1,\ldots,N$ and $q=1,\ldots,p$. Given that, in reality, these are seldom the observed values, we, consecutively, go through all connections where either $g_{ij}=1$ or $X_{ij,1}=x_{1},\ldots,X_{ij,p}=x_{p}$ for $\sum_{q=1}^{p}\,x_{q}\neq 0$ holds and correct for the resulting error. In formulae, we employ the following decomposition:
$\Omega_{ik}^{(s)}(\bm{g},\bm{x},\bm{\xi}^{(s)})~{}$ $\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left\{\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right\}$
$=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left\{\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right\}$
$\hskip 14.22636pt+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left[g_{ij}\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}+g_{ji}\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}\right]$
$\hskip 14.22636pt+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left[(1-g_{ij})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}+(1-g_{ji})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}\right]$
$=\Omega_{ik}^{(s)}\left(\bm{0},\bm{0},\bm{\xi}^{(s)}\right)+\Lambda_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right),$ (17)
@@
$\Lambda_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left[g_{ij}\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}+g_{ji}\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}\right]$ (18)
$+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left[(1-g_{ij})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right\}+(1-g_{ji})\xi_{jl}^{(s)}\left\{\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right\}\right]$
being the error in $\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$ arising from assuming an empty network with all categorical covariates set to zero. Following, we show how both terms, $\Omega_{ik}^{(s)}\left(\bm{0},\bm{0},\bm{\xi}^{(s)}\right)$ and $\Lambda_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\xi}^{(s)}\right)$, can be computed through matrix operations, completing our scalable algorithmic approach.
We, first, assume an empty network with all categorical covariates set to zero, $g_{ij}=x_{q}=0$ for all $i\neq j$ and $q=1,...,p$, to compute $\Omega_{ik}^{(s)}\left(\bm{0},\bm{0},\bm{\xi}^{(s)}\right)$ through matrix multiplications. Observe that
$\Omega_{ik}^{(s)}\left(\bm{0},\bm{0},\bm{\xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
@@
$\bm{A_{0}}^{(s)}~{}\coloneqq~{}\begin{bmatrix}\tau^{(s)}_{1}-\xi_{11}^{(s)}&\tau^{(s)}_{2}-\xi_{12}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{1K}^{(s)}\\ \tau^{(s)}_{1}-\xi_{21}^{(s)}&\tau^{(s)}_{2}-\xi_{22}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{2K}^{(s)}\\ \vdots&\vdots&\ddots&\vdots\\ \tau^{(s)}(1)-\xi_{n1}^{(s)}&\tau^{(s)}_{2}-\xi_{n2}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{nK}^{(s)}\end{bmatrix}$
@@
$\bm{\Pi_{0}}^{(s)}~{}\coloneqq~{}\begin{bmatrix}\log\,\pi_{11}^{(s)}(0,\bm{0})&\log\,\pi_{12}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{1K}^{(s)}(0,\bm{0})\\ \log\,\pi_{21}^{(s)}(0,\bm{0})&\log\,\pi_{22}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{2K}^{(s)}(0,\bm{0})\\ \vdots&\vdots&\ddots&\vdots\\ \log\,\pi_{K1}^{(s)}(0,\bm{0})&\log\,\pi_{K2}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{KK}^{(s)}(0,\bm{0})\end{bmatrix},$
@@
$\bm{A_{0}}^{(s)}\left\{\left(\bm{\Pi_{0}}^{(s)}\right)^{\top}+\bm{\Pi_{0}}^{(s)}\right\}=\left(\Omega_{ik}^{(s)}(\bm{0},\bm{0},\bm{\xi}^{(s)})\right)_{ik}$
holds. Put differently, we are able to write $\Omega_{ik}^{(s)}(\bm{0},\bm{0},\bm{\xi})$ as the $(i,k)$th entry of multiplying $\bm{A_{0}}^{(s)}$ and $\bm{\Pi_{0}}^{(s)}(\bm{z})^{\top}+\bm{\Pi_{0}}^{(s)}(\bm{z})$. The matrix $\pi^{(s)}(1,\bm{0})$ can be computed via sparse matrix operations:
$\bm{\pi}^{(s+1)}(1,\bm{0})=\left\{\left(\bm{\xi}^{(s+1)}\right)^{\top}\bm{g}\circ(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\ \left(\bm{\xi}^{(s+1)}\right)\right\}\oslash$ (19)
$\left\{\left(\bm{\xi}^{(s+1)}\right)^{\top}(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\left(\bm{\xi}^{(s+1)}\right)\right\},$
@@
$\bm{\Pi_{0}}^{(s)}(\bm{z})=\log\left(1-\bm{\pi}^{(s+1)}(d=1,X_{1}=0,\ldots,X_{p}=0)\right)$

### v2

@@ A.1.1 Variational EM algorithm with MM updates
holds with fixed $\bm{\alpha},\bm{\beta},\bm{\eta}$ and $\bm{\Xi}^{(s)}$. Applying these two properties guarantees that maximizing the function $M$ forces the lower bound uphill:
$\begin{array}[]{llllllllll}\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s)})&=&M\left(\bm{\Xi}^{(s)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)&\leq&M\left(\bm{\Xi}^{(s+1)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)\\ &\leq&\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s+1)})\end{array}$
Extending the approach of Vu et al. (2013) to directed networks, the following expression minorizes $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ ::=:= $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})$
@@
In Section B, we detail how sparse matrix multiplication can be exploited to speed up the updates of $\bm{\Xi}^{(s+1)}$.
We run this algorithm until the relative increase of the variational lower bound in below a certain threshold to obtain estimates for $\widehat{\bm{\Xi}}$ and $\widehat{\bm{\eta}}$. We then assign each node to its modal estimated type. Therefore, package $i$ is of type $k$, $\widehat{z}_{ik}=1$, if $\widehat{\bm{\Xi}}_{ik}\geq\widehat{\bm{\Xi}}_{i\ell}$ for all $\ell\neq k$ and for all $i$s.
ONLINE APPENDIX
Appendix B Exploiting sparse matrix operations for updating variational parameters
Next, we detail how the updates of $\mathbf{\xi}$ in Step 1 can be carried out in a scalable manner based on sparse matrix operations. Equation (A.1.1) can be written as a quadratic programming problem in $\mathbf{\xi}$ with the block constraints that $\sum_{k=1}^{K}\xi_{ik}=1$ needs to hold for all $i=1,\ldots,N$. We rearrange the first part of (A.1.1) as follows:
$\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\,\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\,\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{ik}^{2}\,\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\frac{\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})}{2\xi_{ik}^{(s)}}\,\xi_{ik}^{2}$
@@ Appendix B Exploiting sparse matrix operations for updating variational parameters
$\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right).$ (28)
@@
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ $=\sum_{i=1}^{N}\sum_{k=1}^{K}\left(\frac{\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}\right)\xi_{ik}^{2}+\left(\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1\right)\xi_{ik}$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\,\xi_{ik}^{2}+B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\,\xi_{ik}$
@@
$A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\frac{\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}$
@@
$B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1$
@@
To update the estimate of $\bm{\Xi}$, we need to evaluate $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ and $B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$, which, when done naively, is of complexity $O(N^{2}\,K^{2})$. Computing $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ for $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ is the computational bottleneck driving these complexities. It is thus prohibitively difficult for a large number of population members and communities in the network. Note that in our application, this problem is exasperated as we have more than 35,000 nodes and our theoretical result on the adequacy of the estimation procedure assumes that $K$ grows as a function of $N$. To avoid this issue, we show how $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ can be evaluated through a series of sparse matrix multiplications. Thereby, our implementation is fast and light on memory requirements. The underlying idea is that $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ collapses to a simple form if $\mathbf{g}$ is an empty network with all pairwise covariates set to zero, i.e., $g_{ij}=0$ and $x_{q}=0\,\forall\,i,j=1,\ldots,N$ and $q=1,\ldots,p$. Given that, in reality, these are seldom the observed values, we, consecutively, go through all connections where either $g_{ij}=1$ or $X_{ij,1}=x_{1},\ldots,X_{ij,p}=x_{p}$ for $\sum_{q=1}^{p}\,x_{q}\neq 0$ holds and correct for the resulting error. In formulae, this implies the following:
$\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})~{}$ $\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$=\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)+\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right),$ (29)
@@
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$ (30)
$+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
being the error in $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ arising from assuming an empty network with all categorical covariates set to zero. Next, we show how both terms, $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$, can be computed through matrix operations, completing our scalable algorithmic approach.
We, first, assume an empty network where all categorical covariates set to zero, $g_{ij}=x_{q}=0$ for all $i\neq j$ and $q=1,...,p$ and compute $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ via matrix multiplications. Observe that
$\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
@@
$\bm{A_{0}}\left(\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\begin{bmatrix}\tau^{(s)}_{1}-\xi_{11}^{(s)}&\tau^{(s)}_{2}-\xi_{12}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{1K}^{(s)}\\ \tau^{(s)}_{1}-\xi_{21}^{(s)}&\tau^{(s)}_{2}-\xi_{22}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{2K}^{(s)}\\ \vdots&\vdots&\ddots&\vdots\\ \tau^{(s)}(1)-\xi_{n1}^{(s)}&\tau^{(s)}_{2}-\xi_{n2}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{nK}^{(s)}\end{bmatrix}$
@@
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\begin{bmatrix}\log\,\pi_{11}^{(s)}(0,\bm{0})&\log\,\pi_{12}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{1K}^{(s)}(0,\bm{0})\\ \log\,\pi_{21}^{(s)}(0,\bm{0})&\log\,\pi_{22}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{2K}^{(s)}(0,\bm{0})\\ \vdots&\vdots&\ddots&\vdots\\ \log\,\pi_{K1}^{(s)}(0,\bm{0})&\log\,\pi_{K2}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{KK}^{(s)}(0,\bm{0})\end{bmatrix},$
@@
$\bm{A_{0}}\left(\bm{\Xi}^{(s)}\right)\left(\left(\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)\right)^{\top}+\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)\right)=\left(\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)\right)_{ik}$ (31)
holds. Put differently, we are able to write $\Omega_{ik}^{(s)}(\bm{0},\bm{0},\bm{\Xi})$ as the $(i,k)$th entry of the result of (31). To compute $\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)$, we first evaluate the matrix $\pi^{(s)}(1,\bm{0})\in[0,1]^{K\,\times\,K}$ via sparse matrix operations:
$\bm{\pi}^{(s+1)}(1,\bm{0})=\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{g}\circ(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\ \left(\bm{\Xi}^{(s+1)}\right)\right)\oslash$ (32)
$\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\left(\bm{\Xi}^{(s+1)}\right)\right),$
@@
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)=\log\left(1-\bm{\pi}^{(s+1)}(1,\bm{0})\right)$